Lightbulbs of a certain type are advertised as having an average lifetime of 750 hours. The price of these bulbs is very favorable, so a potential customer has decided to go ahead with a purchase arrangement unless it can be conclusively demonstrated that the true average lifetime is smaller than what is advertised. A random sample of 50 bulbs was selected, and the sample mean was found to be 738.44 with a standard deviation of 38.20 and a standard error on the mean of 5.40. What is the p-value for this test (to 4 decimals)?
Answer #1

Solution :

Given that ,

\mu = 750

T = 738.44

\sigma = 38.20

n = 50

The null and alternative hypothesis is ,

H0 :  \mu = 750

Ha : \mu < 750

This is the left tailed test .

Test statistic = z

= (T - \mu ) / \sigma / \sqrt n

= ( 738.44 - 750) / 38.20 / \sqrt 50

= -2.14

The test statistic = -2.14

P - value = P (Z < -2.14 ) = 0.0162

The P-value for the test = 0.0162